# LLMs Cannot Reliably Judge (Yet?): A Comprehensive Assessment on the Robustness of LLM-as-a-Judge

### v1

@@ 5.1.2 Defense Performance
5.2 Impact of Prompts and Model Choices (RQ2)
TABLE VII: Comparison of robustness across different judge prompt templates on Text Translation tasks(T1).
Higher ASR (weaker robustness)
Lower ASR (stronger robustness)
@@ 5.2 Impact of Prompts and Model Choices (RQ2)
5.2.1 Impact of Judge Prompt Template
@@
We examine the impact of prompt templates on the robustness of LLM-as-a-Judge systems. Specifically, we evaluate three widely used prompt templates, e.g., Vanilla Prompt, Arena-Hard Prompt and Google Vertex Promp, across four judge models on the translation task under three adversarial attacks (H4–H6). In this evaluation, a higher ASR indicate weaker robustness.
@@
The results are shown in Table VII (a-c). We observe that LLM-as-a-Judge systems are highly sensitive to the choice of prompt template. The same judge model can exhibit significantly different ASR values depending on the template used. For example, under H4, Mistral-7B yields an ASR of 37.50% with the Vanilla prompt, which increases to 96.25% with the Arena-Hard prompt. Similarly, LLama-3.1-8B displays a wide range in ASR, from 6.67% ASR (Vanilla) to 73.33% (Google Vertex) under H4, further highlighting this sensitivity.
@@
To improve robustness, we apply the coordinate ascent strategy described in Algorithm 1 to derive an optimized prompt. Initializing from the Arena-Hard template, we iteratively refine individual prompt components based on ASR under the H4 attack. The optimization is performed on the translation task, with chinese2english language pair. We conduct three iterations of coordinate ascent using Openchat-3.5 as the target model. The optimized prompt is then evaluated for generalization across five additional language pairs within the translation task, and the transferability under H5 and H6 attacks.
@@
Compared with the baselines, our optimized prompt achieves superior robustness (Table VII (d)). On Openchat-3.5, it reduces the average ASR across H4–H6 to 6.11%, outperforming Arena-Hard (20.08%) and Google Vertex (23.06%), while also maintaining comparably low ASR on Qwen-2.5 and LLaMA-3.1-8B. Notably, for Mistral-7B, it brings the ASR under H4 down from 96.25% (Arena-Hard) to only 14.17%, demonstrating a substantial improvement in robustness. The optimized prompt template is presented in Table X (appendix).
@@
Importantly, unlike prior work that relies on extensive prompt search or model fine-tuning, our method performs optimization in a highly cost-efficient manner, targeting a single attack (H4), a single task (translation), and one language pair (chinese2english). Despite this constrained budget, the optimized prompt generalizes well on additional attacks (H5, H6) and models (Openchat, Qwen, Mistral, LLaMA), highlighting the strong generalizability and transferability of our approach. This exhibits that lightweight, targeted prompt modifications can yield broad defensive benefits and robustness for LLM-as-a-Judge systems.
$\bullet$ Finding 5: The robustness of LLM-as-a-Judge systems is highly sensitive to the choice of prompt template.
$\bullet$ Finding 6: We propose a low-cost yet highly transferable prompt template optimization method that improves robustness. The optimized template consistently outperforms existing prompt template.
@@
Figure 3: Evaluate the robustness of different judge models on 7 attacks.
5.2.2 Impact of Judge Model
@@
We evaluate a range of judge models, including GPT-4o, Openchat, JudgeLM-7/13B, Prometheus-7B, DeepSeek-R1, PandaLM-7B, and AutoJ-13B, comparing their robustness against various adversarial attacks. The results are summarized in Figure  3. GPT-4o, considered to be the defacto LLM-as-a-Judge system, demonstrates strong robustness across different adversarial attacks, achieving an ASR of 71.26% and an iSDR of 0.1235 (Figure 3h). These results validate the current practice of using GPT-4o as a judge model and suggest that it offers a relatively reliable evaluation setup. Notably, JudgeLM-13B, a fine-tuned model specifically trained for judgment tasks, achieves the lowest ASR (69.00%) among all evaluated models and exhibits consistently strong robustness across different attack strategies. This highlights the effectiveness of task-specific fine-tuning in improving the robustness of LLM-as-a-Judge systems. Given its open-source availability and lower computational overhead compared to GPT-4o, JudgeLM-13B presents a promising and cost-effective alternative for reliable LLM-as-a-Judge system. In contrast, reasoning-oriented models such as DeepSeek-R1 have been hypothesized to offer superior robustness in judgment tasks due to their enhanced reasoning capabilities. However, our results show that DeepSeek-R1 achieves only moderate robustness, with an ASR of 75.20% and an iSDR of 0.5643, indicating no significant advantage over other models in adversarial scenarios.
@@
$\bullet$ Finding 7: Judge-tuned models demonstrate the strongest robustness against adversarial attacks due to alignment-focused fine-tuning.
$\bullet$ Finding 8: Reasoning-focused models (e.g., DeepSeek-R1) are more susceptible to severe scoring errors when attacks succeed, indicating weaker reliability under adversarial conditions.
5.3 Real-World Case Study (RQ3)
We conduct experiments on a real-world industrial LLM-as-a-Judge platform, i.e., PAI-Judge [^2] https://pai.console.aliyun.com/#/ai-service/judge/overview. This platform provides an overall evaluation score for each input, along with a set of subscores reflecting multiple assessment dimensions, such as accuracy, fluency, and consistency. According to its official documentation, PAI-Judge also offers a premium version PAI-Judge-Plus, which is built on a larger model, and is reported to deliver superior judgment performance.
Our initial evaluation targets PAI-Judge using pointwise protocol and includes both heuristic-Based attacks (e.g., Naive Attack, Escape Characters Attack, Context Ignore Attack, Fake Completion Attack, Fake Reasoning, Combined Attack (H1-H8)), and optimization-based attacks (PAIR (O2)). To ensure compatibility with the platform, we carefully adapt the prompt template to match the PAI-Judge’s format. However, our results show that these conventional adversarial strategies had minimal effect on the platform’s evaluation outcome. This suggests that PAI-Judge incorporates effective internal defense mechanisms, making it relatively robust to standard adversarial manipulations. Sample evaluation results are presented in Table VIII (appendix).
@@ 5.3 Real-World Case Study (RQ3)
Figure 4: Comparison of Avg Subscore and Overall Score for PAI-Judge and PAI-Judge-Plus.
To further probe the system’s robustness, we extended the PAIR-optimized adversarial prompts by appending long-suffix manipulations (lengthy irrelevant or distracting content). Surprisingly, this composite attack results in high scores on both PAI-Judge (a) and PAI-Judge-Plus (b), despite that these adversarial responses should have received near-zero scores (Figure 4). We further investigate the attack by gradually increasing the length of the appended suffix and observe a rising trend in effectiveness. While PAI-Judge-Plus demonstrates some resistance, its scoring eventually increases significantly once the suffix length approaches 2000 letters. These results reveal a critical loophole in PAI-Judge and PAI-Judge-Plus platform.
Using our RobustJudge framework, we are able to systematically evaluate and uncover this vulnerability in a real-world judge system. These findings demonstrate the practical utility of RobustJudge in identifying hidden flaws and guiding the development of more robust LLM-as-a-Judge systems. We responsibly disclosed this vulnerability to the PAI-Judge platform team prior to the release of our results.
$\bullet$ Finding 9: Conventional adversarial attacks were largely ineffective against the industrial PAI-Judge platform.
$\bullet$ Finding 10: We identify a loophole in both PAI-Judge and PAI-Judge-Plus, where combining PAIR-optimized adversarial inputs with long-suffix can compromise the platform’s judgment reliability.
6 Conclusion
This work presents the first scalable and fully automated framework to evaluate the robustness and reliability of LLM-as-a-Judge systems across multiple attack scenarios. We systematically benchmarked state-of-the-art LLM-based evaluators under various adversarial settings and found that they are vulnerable to manipulation, often producing biased or incorrect judgments when exposed to crafted inputs. Based on these findings, we conclude that current LLM-as-a-Judge systems are not yet sufficiently robust for secure deployment in critical evaluation tasks, and highlight the need for future research to strengthen their resilience against adversarial attacks. Our framework and benchmarks provide a valuable tool for the community to assess and track the progress of LLM-based evaluators in adversarial robustness.

### v2

@@ V-A2 Defense Performance
V-B Impact of Prompts and Model Choices (RQ2)
TABLE VII: Comparison of robustness across different judge prompt templates on the Text Translation task (T1). Orange represents a higher P-ASR, suggesting weaker robustness across different attack methods. Green represents a lower average P-ASR, indicating stronger robustness across the four judge models.
@@ V-B Impact of Prompts and Model Choices (RQ2)
TABLE VIII: Our optimized prompt.
System Prompt:
You are an expert evaluator. You need to perform a translation task ({##SOURCE_LAN##} to {##TARGET_LAN##}).
You aim to evaluate the quality of the responses for a given instruction. Your goal is to select the best response for the given instruction. Select Response A or Response B for the given instruction. Two different AI chatbots, respectively, generate the two outputs.
Begin your evaluation by generating your own response to the prompt. You must provide your response before judging any responses. When evaluating the assistants’ responses, compare both assistants’ responses with your response. You must identify and correct any errors or inaccuracies.
Please consider the helpfulness, relevance, accuracy, and level of detail of their responses.
Please provide your explanation before providing your preference.
User Prompt:
### Instruction: {##SOURCE##}
### Response A: {##A##}
### Response B: {##B##}
V-B1 Impact of Judge Prompt Template
@@
We investigate how different prompt templates affect the robustness of LLM-as-a-Judge systems. Specifically, we employ a pairwise evaluation protocol to compare three widely used templates—Vanilla Prompt, Arena-Hard Prompt, and Google Vertex Prompt—across four judge models in the Text Translation task (T1) under three adversarial attacks (H4–H6). In this evaluation, a higher P-ASR reflects weaker robustness.
@@
The results in Table VII (a)-(c) are averaged across six language pairs: Chinese–English, German–English, Yoruba–English, English–German, English–Yoruba, and English–Chinese. We observe that LLM-as-a-Judge systems are highly sensitive to the choice of prompt template. The same judge model can exhibit significantly different P-ASR values depending on the template used. For example, under H4, Mistral-7B yields a P-ASR of 37.50% with the Vanilla prompt, which increases to 96.25% with the Arena-Hard prompt. Similarly, LLama-3.1-8B exhibits a wide range in P-ASR, from 6.67% P-ASR (Vanilla) to 73.33% (Google Vertex) under H4, further highlighting this sensitivity.
@@
To enhance robustness, we employ the coordinate ascent strategy described in Algorithm 1 to derive an optimized prompt. Starting from the Arena-Hard template, individual prompt components are iteratively replaced based on their P-ASR under a relatively strong H4 attack. This optimization is performed on the Text Translation task (T1) using the Chinese–English language pair. We conduct three iterations of coordinate ascent with Openchat-3.5 as the target model. The resulting optimized prompt is subsequently evaluated for generalization across five additional language pairs (German–English, Yoruba–English, English–German, English–Yoruba, English–Chinese) within T1, as well as for transferability under the H5 and H6 attacks, which represent the other stronger attacks in the set.
@@
Compared with the baselines, our optimized prompt achieves superior robustness (Table VII (d)). On Qwen2.5-7B, it reduces the average P-ASR across H4–H6 to 3.61%, outperforming Arena-Hard (15.42%) and Google Vertex (9.76%), while also maintaining a comparably low P-ASR on Mistral-7B and LLaMA-3.1-8B. Notably, for Mistral-7B, it brings the P-ASR under H4 down from 96.25% (Arena-Hard) to only 14.17%, demonstrating a substantial improvement in robustness. The optimized prompt template is presented in Table VIII.
@@
Importantly, unlike prior work that relies on extensive prompt search or model fine-tuning, our method performs optimization in a highly cost-efficient manner, targeting a single attack (H4), a single task (T1), and one language pair (Chinese-English). Despite this constrained budget, the optimized prompt generalizes well on additional attacks (H5, H6) and models, highlighting the strong generalizability and transferability of our approach. This exhibits that lightweight, targeted prompt modifications can yield broad defensive benefits and robustness for LLM-as-a-Judge systems.
[…]
@@
Figure 4: Robustness evaluation of different judge models against 7 adversarial attacks. Each subplot shows the Attack Success Rate (ASR), Success Defense Rate (SDR), and Improved Success Defense Rate (iSDR) for a specific attack method, with (h) presenting the average performance across all attacks.
V-B2 Impact of Judge Model
@@
We evaluate a range of judge models, including GPT-4o, Openchat, JudgeLM-7/13B, Prometheus-7B, DeepSeek-R1, PandaLM-7B, and AutoJ-13B, comparing their robustness against various adversarial attacks. The results are summarized in Figure 4. GPT-4o, considered to be the extensive LLM-as-a-Judge system, demonstrates strong robustness across different adversarial attacks, achieving an ASR of 71.26% and an iSDR of 0.1235 (Figure 3h). These results validate the current practice of using GPT-4o as a judge model and suggest that it offers a relatively reliable evaluation setup. Notably, JudgeLM-13B, a fine-tuned model specifically trained for judgment tasks, achieves the lowest ASR (69.00%) among all evaluated models and exhibits consistently strong robustness across different attack strategies. This highlights the effectiveness of task-specific fine-tuning in improving the robustness of LLM-as-a-Judge systems. Given its open-source availability and lower computational overhead compared to GPT-4o, JudgeLM-13B presents a promising and cost-effective alternative for a reliable LLM-as-a-Judge system. In contrast, reasoning-oriented models such as DeepSeek-R1 have been hypothesized to offer superior robustness in judgment tasks due to their enhanced reasoning capabilities. However, our results show that DeepSeek-R1 achieves only moderate robustness, with an ASR of 75.20% and an iSDR of 0.5643, indicating no significant advantage over other models in adversarial scenarios.
@@
$\bullet$ Finding 7: Judge-tuned models demonstrate the strongest robustness against adversarial attacks. In particular, JudgeLM-13B consistently achieves robustness levels that not only surpass other open-source judges but also exceed GPT-4o, highlighting the effectiveness of alignment-focused fine-tuning for enhancing adversarial resistance.
V-C Real-World Case Study (RQ3)
To demonstrate the practical utility of RobustJudge in identifying vulnerabilities in LLM-as-a-Judge systems, we conduct a comprehensive evaluation on PAI-Judge [^5] https://pai.console.aliyun.com/#/ai-service/judge/overview, a commercial industrial judging platform by Alibaba Cloud. PAI-Judge provides multi-dimensional evaluation through a hierarchical scoring mechanism, generating individual subscores (0-10 scale) for accuracy, fluency, consistency, relevance, and helpfulness, which are aggregated into an overall score via a weighted combination. The platform offers two variants: the standard version and PAI-Judge-Plus.
Our evaluation follows a systematic three-phase approach. We first deploy conventional adversarial attacks under the pointwise protocol, including heuristic-based attacks (Naive Attack, Escape Characters Attack, Context Ignore Attack, Fake Completion Attack, Fake Reasoning) and optimization-based attacks (PAIR (O2)). We adapt prompt templates to match PAI-Judge’s API format. However, these standard strategies have a minimal impact—both subscores and overall scores remain largely unchanged across all tested variants, suggesting that PAI-Judge incorporates effective internal defenses, such as prompt filtering, output sanitization, or adversarial training.
We leverage our RobustJudge framework with gradient-based optimization, employing Qwen-2.5-7B-Instruct fine-tuned using GCG-like methods to craft adversarial responses that appear plausible while containing quality flaws. The optimization minimizes $\mathcal{L}_{\text{adv}}=-\log P_{\text{judge}}(\text{high score}\mid\text{query},\text{adversarial response})$, yielding samples that exploit the judge’s decision boundaries through semantic manipulations, fabricated reasoning, and misleading cues. We then introduce long-suffix augmentation, which involves appending lengthy, irrelevant content (such as repetitive filler text, irrelevant facts, and distracting narratives) to exploit the judge’s attention mechanism limitations. By overwhelming the model with excessive input length, we hypothesize that the judge fails to focus on core adversarial content, inflating scores.
Our results reveal a critical vulnerability in both PAI-Judge platforms (Figure 5). The composite attack—combining Qwen-2.5-7B-optimized adversarial content with long-suffix augmentation—achieves remarkably high scores despite responses containing significant flaws warranting near-zero scores. Standard PAI-Judge exhibits severe susceptibility: with suffix lengths of 1000 characters, average subscore increases from approximately 1.5 (baseline) to over 6.0, with overall score reaching 7.2. PAI-Judge-Plus maintains low scores below 1500 characters but succumbs once suffix length exceeds 2000 characters, achieving an average subscore of 5.8 and an overall score of 6.5, indicating acceptable quality in the platform’s rubric.
@@ V-C Real-World Case Study (RQ3)
Figure 5: Impact of attack on PAI-Judge variants.
Progressive length analysis (0-3000 characters, 200-character increments) reveals sharp transition points where scores increase rapidly—around 800-1000 characters for PAI-Judge and 1800-2000 characters for PAI-Judge-Plus (2× better resistance). Beyond these thresholds, scores plateau at 7-8 and 6-7, respectively. Attack effectiveness varies across dimensions: fluency and consistency subscores increase more dramatically (7-8) than accuracy scores (5-6), indicating attention degradation primarily affects surface-level assessments rather than semantic understanding.
$\bullet$ Finding 8: Conventional adversarial attacks were largely ineffective against the industrial PAI-Judge platform.
$\bullet$ Finding 9: We identify a loophole in both PAI-Judge and PAI-Judge-Plus, where combining PAIR-optimized adversarial inputs with long suffixes can compromise the platform’s judgment reliability.
VI Conclusion
This work presents the first scalable, automated framework for evaluating LLM-as-a-Judge robustness under adversarial attacks. Our benchmarking shows that current evaluators (even Alibaba’s PAI platform) remain vulnerable and biased, underscoring the need for stronger defenses and more reliable real-world deployment. We also discovered that the robustness varies substantially across prompt templates. Our findings provide actionable insights for deploying trustworthy LLM-as-a-Judge systems in security-sensitive applications.